0 0
Add a comment Improve this question Transcribed image text
Answer #1

A) is correct.

\int_{0}^{4}\int_{\sqrt{x}}^{2}(\frac{x^{2}}{y^{7}+1})dy dx=\int_{0}^{2}\int_{0}^{y^{2}}(\frac{x^{2}}{y^{7}+1})dx dy

\int_{0}^{2}\int_{0}^{y^{2}}(\frac{x^{2}}{y^{7}+1})dx dy=\int_{0}^{2}(\frac{x^{3}}{3(y^{7}+1)})|_{0}^{y^{2}} dy=\int_{0}^{2}(\frac{y^{6}}{3(y^{7}+1)}) dy=log(129)/21=0.2314